OpenStudy (anonymous):

Simplify the radical. sqrt 125 would it be 5 sqrt 25 ?

OpenStudy (anonymous):

no it ll b 5 sqrt 5

OpenStudy (anonymous):

sqrt(125) sqrt(5*25) sqrt(5*5^2) 5sqrt(5)

OpenStudy (callisto):

\[\sqrt{125} = \sqrt{5\times 5 \times 5} =\sqrt{5^2 \times 5}=5\sqrt{5}\]

OpenStudy (anonymous):

@AlexisDelgado i can help u in ol dis kinda questoin keep asking ..